OpenStudy (surryyy):

Find f(5) if f(x) = 2 (3x-5/x-1) 5 20 4 2

OpenStudy (anonymous):

what it means is for you to put replace x as 5

imqwerty (imqwerty):

f(5) is the value which u get on putting x=5 in the expression :)

OpenStudy (jiteshmeghwal9):

Put 5 on the place of \(x\) \[f(5)=2\left( \frac{3(5)-5}{5-1} \right)\]

OpenStudy (anonymous):

\[F(5)=2((3(5)-5)/(5)-1)\]

OpenStudy (anonymous):

answer 4

OpenStudy (surryyy):

sorry was lagging. but i understand now. tysm.

imqwerty (imqwerty):

( ͡° ͜ʖ ͡°)